Evaluate the integral "\displaystyle{\int_{0}^{1}\frac{x}{x+1}dx}"

  1. "1-\ln{ 2}"
  2. "\ln{2}"
  3. 0
  4. "-\ln{2}"





1
Expert's answer
2021-09-06T16:33:50-0400

"\displaystyle\int_{0}^{1}\dfrac{x}{x+1}dx=\displaystyle\int_{0}^{1}\dfrac{x+1-1}{x+1}dx"

"=\displaystyle\int_{0}^{1}dx-\displaystyle\int_{0}^{1}\dfrac{1}{x+1}dx"

"=[x]\begin{matrix}
 1 \\
 0
\end{matrix}-\ln(|x+1|)\begin{matrix}
 1 \\
 0
\end{matrix}"

"=1-0-(\ln(|1+1|)-\ln(|0+1|))"

"=1-\ln2"



1. "1-\ln2"